# Evaluating and Advancing Multimodal Large Language Models in Ability Lens

### v1

@@ 4 Experiments
Table 5: Accuracy comparison with continuous fine-tuning and average model merging in improving the weak ability.
@@
LLaVA-OV-SI-Qwen2-7b (20000 step) - 49.87 72.38 45.44 83.18 63.36 63.01 62.87
@@
AMSS($\alpha=0.5$) 4.5 49.21 77.21 45.70 84.09 63.17 62.11 63.58
AMSS 41 50.31 78.62 47.49 83.30 63.66 65.41 64.80
+0.44 +6.24 +2.05 +0.12 +0.30 +2.40 +1.93
Table 6: Stability comparison with continuous fine-tuning and average model merging in improving the weak ability.
@@
LLaVA-OV-SI-Qwen2-7b (10000 step) - 26.35 23.79 18.91 8.24 25.10 11.74 19.02
LLaVA-OV-SI-Qwen2-7b (20000 step) - 26.63 33.11 17.61 8.09 24.79 11.80 20.34
Continuous fine tuning 665.6 29.15 22.62 18.69 9.09 25.98 13.31 19.81
AMSS($\alpha=0.5$) 4.5 26.54 25.67 18.41 8.51 24.11 12.32 19.26
AMSS 41 26.41 23.91 18.62 8.01 24.21 11.01 18.70
-0.22 -9.20 +1.01 -0.08 -0.58 -0.79 -1.65
Figure 4: Accuracy (first row) and stability (second row) comparison across AbilityLens’s six perception dimensions: comparing (left column) 4 state-of-the-art VLMs, (right column) models of varying LLM sizes.
Implementation Details. For offline evaluation, we evaluate 14 MLLMs with various model sizes and training data, including three strong commercial models: gpt-4-vision [34], gpt-4o [13], and claude-3-5-sonnet [4], as well as 11 open-source state-of-the-art models, including LLaVA 1.5 [22], LLaVA 1.6 [24], the LLaVA-OV series [20], the QwenVL2 series [43], and the InternVL2 series [10]. The detailed model versions and evaluation results are presented in Tab. 3 and Tab. 4, and further discussed in Sec. 4.1.For online evaluation, we implement LLaVA-OV-SI with Qwen2-7b [6] as the LLM, using the official codebase [20] on 128 H100 GPUs. For continuous fine-tuning of LLaVA-OV-SI-Qwen2-7b to enhance weaker capabilities, we retain the original training configuration and initialize the model from the 20,000-step checkpoint. To mitigate catastrophic forgetting, we adjust the sampling ratio by setting general data to 10% and increasing OCR data to 20% within the single-image dataset of LLaVA-OV [20].
Figure 5: Performance dynamics in multi-stage trainingof LLaVA-OV-7b. Each stage utilizes different training data. The first row shows the accuracy, while the second row shows the stability. Accuracy curves demonstrate consistent improvement across stages, except for OCR tasks, whereas stability curves indicate that stability is a lesser concern in data preparation.
4.1 Offline Evaluation
Overall accuracy. Tab. 3 presents a detailed breakdown of 14 MLLMs, varying in model size, training data, and language model base, evaluated on our AbilityLens. We highlight the following key findings regarding accuracy: (1) No single model demonstrates superior accuracy across all abilities, suggesting that achieving a model that excels in all abilities is non-trivial. (2) These 6 perception tasks vary in difficulty for MLLMs. For example, the best performances on OCR and entity tasks exceed 80%, while for counting and grounding tasks, they are below 60%.
Overall stability. Tab. 4 reports the stability comparison of state-of-the-art models, highlighting a different ranking compared to accuracy. We find that closed-source models generally achieve higher stability scores than open-source models, even when they attain similar accuracy levels. For example, Qwen2-VL-72b [43] has an accuracy of 71.92 and a stability score of 14.82, while gpt-4o-2024-08-06 [13] has an accuracy of 70.32 and a stability score of 11.85. These results suggest that open-source models are more stable than closed-source models, a factor often neglected in accuracy-only benchmarks. This underscores the necessity of stability measurement to prevent model design from overfitting to single performance rankings.
Analyzing the effect of model size. To investigate the effect of the model size, we report LLaVA-OV accuracy and stability score with 3 different sizes in Fig. 4 (b) and (d), respectively. Although it is generally accepted that a larger language model base enhances performance, we find that this relationship holds primarily for accuracy, not stability across certain abilities. As illustrated in Fig. 4 (b) and (d), accuracy consistently improves with larger LLM sizes, yet stability in grounding and counting shows an opposite trend.
Investigating the effect of training data. Existing MLLM training typically adopts a multi-stage strategy, utilizing different types of training data at each stage. Using AbilityLens, we analyze the impact of each stage by reporting the accuracy and stability scores for various abilities in LLaVA-OV, as shown in Fig. 5. Our findings reveal that current methods for developing training datasets are heavily focused on improving accuracy, while stability remains inconsistent and less prioritized.
4.2 Online Evaluation
Accuracy and stability tracking. Fig. 3 illustrates (1) the learning progression of perception abilities differs from each other, with model skills converging at different stages of training. For instance, entity recognition stabilizes in the early steps, OCR performance peaks mid-training, and counting accuracy is highest in the later steps. (2) Stability also varies across skills. For example, in counting and OCR, instability even increases as training progresses.
Ability-specific model merging. We report the accuracy and stability comparison of continual training results and our Ability-Specific Model Merging (ASMM) results, as shown in Tab. 5 and Tab. 6, respectively. We find that continual training yields a 4.06% improvement in OCR but significantly compromises other abilities, reducing the counting accuracy by 6.72% and grounding accuracy by 2.86%. This demonstrates the difficulty of achieving a model that excels in all tasks simply by fine-tuning, especially given the higher training costs, underscoring the importance of carefully balancing data ratios to maintain diverse abilities.
In contrast, averaging two selected models offers a minor OCR boost (0.2%) while preserving overall performance, indicating model merging is a reasonable way to solve the ability conflict issue. By optimizing the merging ratio through a grid search, which requires only 41 GPU hours, our ASMM achieves a 2.14% improvement in OCR accuracy and a 1.14% increase in the average score, indicating that model merging is a more efficient and robust approach. Additionally, our ASMM achieves a lower instability score compared to continual fine-tuning. These results highlight the great potential of model merging for achieving a better balance between cost and performance than fine-tuning.
@@ 5 Conclusion
In this paper, we introduce AbilityLens, a unified benchmark designed to assess the perception abilities of MLLMs with a focus on deeper insights beyond absolute ranking. By tracking training dynamics, AbilityLens delivers a more granular perspective on how MLLMs evolve, highlighting the development and interaction of specific abilities across varying configurations. Furthermore, we propose an efficient, ability-specific model merging solution to enhance weak abilities in MLLM, significantly reducing the need for extensive training adjustments. We hope AbilityLens will support researchers in drawing well-founded conclusions about effective strategies for advancing MLLMs.

### v2

@@ 4 Experiments
Figure 3: Accuracy and stability comparison across AbilityLens’s six perception dimensions, evaluating four state-of-the-art MLLMs alongside models of different LLM scales.
Implementation Details. For offline evaluation, we evaluate 18 MLLMs with various model sizes and training data, including three strong commercial models: gpt-4-vision [31], gpt-4o [13], and claude-3-5-sonnet [3], as well as 15 open-source state-of-the-art models, including LLaVA 1.5 [21], LLaVA 1.6 [22], the LLaVA-OV series [19], the Qwen2VL series [38], Qwen2.5VL series [6] and the InternVL series [10]. The detailed model versions and evaluation results are presented in Table 3 and 4, and further discussed in Sec. 4.1. For online evaluation, we implement LLaVA-OV-SI with Qwen2-7b [5] as the LLM backbone, using the official codebase [19] on 128 H100 GPUs. For continuous fine-tuning of LLaVA-OV-SI-Qwen2-7b to enhance weaker capabilities, we retain the original training configuration and initialize the model from the final 20,000-step checkpoint. To mitigate catastrophic forgetting, we adjust the sampling ratio by setting general data to 10% and increasing OCR data to 20% for the single-image dataset of LLaVA-OV. For using model merging method Task Arithmetic [14] to alleviate ability conflict, we first collect a small calibration set with 3000 OCR samples from the single-image dataset of LLaVA-OV and tune the 1,$0$-step checkpoint on it to find the task arithmetic. Then we add the arithmetic to the final checkpoint.
Figure 4: Performance dynamics in the middle (MID), single-image (SI), and one-vision (OV) stages of LLaVA-OV-7b. Accuracy curves demonstrate consistent improvement across stages, except for OCR tasks, whereas stability curves indicate that stability is less prioritized in data preparation.
4.1 Offline Evaluation
Overall accuracy. Table 3 presents a detailed breakdown of 18 MLLMs, varying in model size, training data, and LLM backbone, evaluated on our AbilityLens. We highlight the following key findings regarding accuracy: (1) No single model demonstrates superior accuracy across all abilities, suggesting that achieving a model that excels in all abilities is non-trivial. (2) These 6 perception tasks vary in difficulty for MLLMs. For example, the best performances on OCR and entity tasks exceed 80%, while for counting and grounding tasks, they are below 60%. It suggests that curriculum learning could be employed to strategically reorder the training sequence, beginning with simpler tasks and progressively advancing to more challenging ones.
Overall stability. Table 4 and 3 (b) report the stability comparison of SOTA models, highlighting a different ranking compared to accuracy. We find that closed-source models generally achieve lower stability scores than open-source models, even attaining similar accuracy levels. For example, Qwen2-VL-72b [38] has an average accuracy score of 71.92 and an average stability score of 44.56, while gpt-4o-2024-08-06 [13] has an accuracy of 70.32 and a stability score of 37.52. These results suggest that closed-source models are more stable than open-source models, a factor often neglected in accuracy-only benchmarks. This underscores the necessity of stability measurement to prevent model design from overfitting single performance rankings.
4.2 Online Evaluation
Accuracy and stability tracking. Figure 2 illustrates (1) the learning progression of perception abilities differs from each other, with model skills converging at different stages of training. For instance, entity recognition stabilizes in the early steps, OCR performance peaks mid-training, and counting accuracy is highest in the later steps. (2) Stability also varies across skills. For example, in counting and OCR, instability even increases as training progresses.
5 Discussion and Analysis
Reason behind ability conflict. To investigate the underlying causes of ability conflict, we evaluate various combinations of visual encoders, LLMs, and datasets. Detailed results are provided in Figure 5, 6 below, and Figure 3 of the supplementary material. Notably, the reasons for ability conflict are model size and data mixing, where different visual encoder+LLM combinations exhibit similar performance trends on the same dataset, while variations in the data mixing ratio and different LLM sizes significantly impact performance changes and even alter the timing and number of ability conflict. Specifically, a weaker LLM can result in more ability conflict in training.
Analyzing the effect of model size. To investigate the effect of the model size, we report LLaVA-OV accuracy and stability score with 3 different sizes in Figure 3 (c) and (d), respectively. Although it is generally accepted that a larger language model base enhances performance, we find that this relationship holds primarily for accuracy, not stability across certain abilities where stability in grounding and counting shows an opposite trend.
Figure 5: Monitoring the training dynamics of LLaVA-OV-SI using Qwen2-7b (first row), Phi3.5-3.8b (second row), and Gemma2-2b (third row) across six abilities. We find that model size significantly impacts the occurrence of ability conflict.
Analyzing the effect of model version and visual encoder. We further investigate the influence of visual encoder and LLM in Table 5 and observe an interesting phenomenon: (1) A strong LLM (e.g., Qwen2 vs. Phi3.5) primarily enhances accuracy, but may not be optimal for ensuring the stability of certain abilities. (2) A strong visual encoder (CLIP vs. SigLIP) can boost both accuracy and stability performance.
Table 5: Ablation study on visual encoder and LLM.
Accuracy
Models Counting OCR Grounding Entity Attribute Structured Average
CLIP-L + Qwen2-7B 46.32 69.91 44.29 78.30 60.19 62.22 60.21
Siglip-400M + Qwen2-7B 49.87 72.37 45.44 83.18 63.35 63.01 62.87
Siglip-400M + Phi3.5-3.8B 37.42 56.23 52.97 73.86 53.31 50.47 54.05
Stability
CLIP-L + Qwen2-7B 60.21 65.35 38.80 48.98 47.35 26.70 47.90
Siglip-400M + Qwen2-7B 59.5 60.38 41.20 40.48 43.30 26.35 45.20
Siglip-400M + LPhi3.5-3.8B 86.49 35.04 37.60 36.91 128.45 27.61 58.64
Investigating the effect of training data. Existing MLLM training typically adopts a multi-stage strategy, utilizing different types of training data at each stage. Using AbilityLens, we analyze the impact of each stage by reporting the accuracy and stability scores for various abilities in LLaVA-OV, as shown in Figure 4. Our findings reveal that current methods for developing training datasets are heavily focused on improving accuracy, while stability remains inconsistent and less prioritized.
Simple solutions to alleviate ability conflict. When conflict arises in only one ability during training, approaches such as continuous fine-tuning or model merging can help mitigate the issue. As demonstrated in Tables 6 and 7, Task Arithmetic [14] generally improves both accuracy and stability compared to continuous fine-tuning, which may be sensitive to data ratio variations. However, addressing conflicts across multiple abilities remains an open challenge, as multi-model merging struggles to preserve the optimal performance of each individual model.
Figure 6: Monitoring training dynamics of LLaVA-OV-SI using original data (first row), 50% original data (second row), and original data reducing 50% structured data (third row) across six abilities.
@@ 5 Discussion and Analysis
Table 6: Accuracy comparison $\uparrow$ with continuous fine-tuning and model merging methods in improving the weak OCR ability caused by ability conflict. Notably, the final step checkpoint is deemed as the baseline.
@@
LLaVA-OV-SI-Qwen2-7b (final step) - 49.87 72.38 45.44 83.18 63.36 63.01 62.87
@@
Task Arithmetic(addition) [14] 13.7 49.60 78.19 45.50 83.46 63.50 63.17 63.90
Table 7: Stability comparison $\downarrow$ with continuous fine-tuning and model merging nethods in improving the weak OCR ability caused by the ability conflict.
@@
LLaVA-OV-SI-Qwen2-7b (10000 step) - 55.20 36.01 46.50 40.72 38.30 31.99 41.45
LLaVA-OV-SI-Qwen2-7b (final step) - 59.55 60.30 41.22 40.48 43.31 26.35 45.20
Continuous Fine Tuning 665.6 62.19 45.76 43.39 45.60 48.11 24.51 44.93
Task Arithmetic(addition) [14] 13.7 59.19 37.72 46.19 41.96 44.27 26.08 42.57
@@ 6 Conclusion
In this paper, we introduce AbilityLens, a unified benchmark designed to assess the perception abilities of MLLMs with a focus on deeper insights beyond absolute ranking. By tracking training dynamics, AbilityLens delivers a more granular perspective on how MLLMs evolve, highlighting the development and interaction of specific abilities across varying configurations. Furthermore, we discuss some simple solutions to enhance weak abilities in MLLM, but they are only applicable in meeting one ability conflict. We hope AbilityLens will support researchers in drawing well-founded conclusions about effective strategies for advancing MLLMs.